OpenStudy (anonymous):

how many grams in n2h2 are in 1.38 mol of a compound

OpenStudy (photon336):

Set up a proportion: you need to figure out the molar mass of N2H2, in 1 mol 1mol/30grams = 1.38mol/x Cross multiply 30grams*1.38 = grams

OpenStudy (photon336):

@dotuechere

OpenStudy (whpalmer4):

Slightly different route to same answer: \[1 \text{ mole N}_2\text{H}_2 = 30.0293 \text{ g N}_2\text{H}_2 \](look up atomic masses of H and N on your periodic table and add the appropriate multiples of each to get the molar mass of the compound) \[1.38 \text{ mole N}_2\text{H}_2 * \frac{30.0293 \text{ g N}_2\text{H}_2 }{1 \text{ mole N}_2\text{H}_2 } =1.38 \cancel{\text{ mole N}_2\text{H}_2} * \frac{30.0293\text{ g N}_2\text{H}_2 }{1 \cancel{\text{ mole N}_2\text{H}_2 }} = \]\[=1.38*30.0293\text{ g N}_2\text{H}_2 \]